# MBSS-T1: Model-Based Self-Supervised Motion Correction for Robust Cardiac T1 Mapping

### v2

@@ Abstract
T1 mapping is a valuable quantitative MRI technique for diagnosing diffuse myocardial diseases. Traditional methods, relying on breath-hold sequences and echo triggering, face challenges with patient compliance and arrhythmias, limiting their effectiveness. Image registration can enable motion-robust T1 mapping, but inherent intensity differences between time points pose a challenge. We introduce MBSS-T1, a self-supervised model for motion correction in cardiac T1 mapping, constrained by physical and anatomical principles. The physical constraints ensure expected signal decay behavior, while the anatomical constraints maintain realistic deformations. The unique combination of these constraints ensures accurate T1 mapping along the longitudinal relaxation axis. MBSS-T1 outperformed baseline deep-learning-based image registration approaches in a 5-fold experiment on a public dataset of 210 patients (STONE sequence) and an internal dataset of 19 patients (MOLLI sequence). MBSS-T1 excelled in model fitting quality (: 0.975 vs. 0.941, 0.946), anatomical alignment (Dice score: 0.89 vs. 0.84, 0.88), and expert visual quality assessment for the presence of visible motion artifacts (4.33 vs. 3.38, 3.66). MBSS-T1 has the potential to enable motion-robust T1 mapping for a broader range of patients, overcoming challenges such as arrhythmias and suboptimal compliance, and allowing for free-breathing T1 mapping without requiring large training datasets. Our code will be publicly available upon acceptance.
1 Introduction
@@
Quantitative T1 mapping is an advanced MRI technique designed to precisely measure intrinsic longitudinal relaxation time in myocardial tissue [27]. This method is increasingly recognized as essential for assessing diffuse myocardial diseases such as inflammation, fibrosis, hypertrophy, and infiltration [27, 29]. “Native” T1 mapping, performed without a paramagnetic contrast agent, effectively detects myocardial conditions, including edema, iron overload, infarcts, and scarring [25]. Furthermore, by allowing for the objective quantification of tissue characteristics, T1 mapping enables longitudinal tracking of changes in the myocardium. This feature is precious for clinical trials, where monitoring the efficacy of interventions over time is critical [25].
@@
Figure 1 presents the process of generating accurate T1 maps. It begins with capturing a series of images at different time points. Each image captures the myocardium at a distinct point during the longitudinal relaxation. These images are then analyzed to calculate T1 relaxation times through signal decay model fitting, resulting in detailed T1 maps that reflect the myocardial tissue properties. However, heart motion, respiration, and spontaneous patient movements can cause significant distortions in T1 maps, affecting their reliability and clinical usefulness and potentially leading to incorrect diagnoses [28]. Echo-triggering is a well-established method for accounting for cardiac motion. Breath-hold sequences, such as the Modified Look-Locker Inversion Recovery (MOLLI) sequence and its variants [24], are often used to minimize respiratory motion artifacts. However, requiring patients to hold their breath presents practical challenges, as not all patients can fully adhere to these instructions. This breath-hold acquisition further limits the number of images that can be captured and makes T1 mapping unsuitable for patients who cannot tolerate breath-holding [24]. Additionally, inadequate echo-triggering due to cardiac arrhythmias can lead to unreliable T1 maps, compromising diagnostic accuracy.
@@
Reducing motion artifacts and enabling free-breathing cardiac T1 mapping can be accomplished through image registration to align scans taken at different time points. However, achieving precise registration is challenging due to inherent complexities in the image data. These complexities include contrast inversion, partial volume effects, and signal nulling for images captured near the zero crossing of the T1 relaxation curve (Fig. 1).
@@
Despite advancements in classical and deep learning-based methods for motion correction and T1 mapping, significant limitations persist. Classical model-based approaches, while effective, are often slow due to their iterative nature [28, 31]. On the other hand, deep-learning-based methods rely heavily on extensively annotated datasets, which are not readily available for all scanners and protocols [8, 32]. Furthermore, these methods typically use either pairwise loss functions based on metrics such as mutual information to guide the registration process [1] or groupwise approaches such as the principal component analysis (PCA), which ignore the signal decay relaxation model [34]. This omission may result in physically unlikely deformations due to changes in image contrast. These limitations hinder the practical applicability and robustness of current methods.
@@
This work presents MBSS-T1, a self-supervised, physically, and anatomically constrained deep learning model for simultaneous motion correction and T1 mapping from either breath-hold or free-breathing acquisitions. Our method integrates physical and anatomical insights to correct motion in T1 mapping. The main innovation lies in incorporating the signal decay model within the network’s architecture in a self-supervised framework while utilizing a pretrained segmentation network to produce segmentations that guide the registration and ensure the anatomical correctness of the deformed images. This approach enables the generation of corrected T1 maps across various imaging protocols, eliminating the need for protocol-specific data collection, manual annotations, and extensive training. This work extends our previous supervised PCMC-T1 model [13], which required a large annotated dataset and extensive training, thereby limiting its generalization capacity across different acquisition settings, such as sequence type and the number of frames acquired.
@@ 2 Related Work
Previous works aimed to address the challenge of motion-robust cardiac T1 mapping can be divided into classical approaches and more recent deep-learning-based approaches. Classical approaches have primarily relied on model-based techniques and iterative algorithms to mitigate the effects of intensity variations between the different frames. More recently, deep learning-based methods have emerged, leveraging large datasets and neural network architectures to enhance motion correction and T1 mapping accuracy. In the following, we review these works.
@@
Zhang et al. addressed intensity variations in their pairwise non-rigid elastic registration framework for T1 mapping by employing the normalized gradient fields difference as their similarity metric [33]. Similarly, Roujol et al. presented an adaptive registration method for varying contrast-weighted images. This approach uses a local non-rigid registration framework based on optical flow to simultaneously estimate motion fields and intensity variations, effectively managing significant contrast differences. Additionally, the method incorporates a regularization term to constrain the deformation field through automatic feature tracking [23].
@@
Guyader et al. introduced a groupwise image registration technique for quantitative MRI that utilizes a dissimilarity metric derived from an approximated form of total correlation. This method addresses the challenge of aligning multiple images simultaneously without relying on a reference image, thereby reducing registration bias and improving the consistency of quantitative measurements [10]. Tao et al. introduced a robust motion correction method for myocardial T1 and extracellular volume mapping using a PCA-based groupwise image registration technique, which registers all MOLLI frames simultaneously with a total-correlation-based metric [26]. Similarly, Huizinga et al. developed a PCA-based groupwise image registration technique for quantitative MRI (qMRI), including T1 mapping, to correct misalignment without a reference image, thus avoiding registration bias. Their method registers all images to a mean space by minimizing a PCA-derived cost function, assuming that a low-dimensional signal model can represent intensity changes in qMRI without needing details of the specific acquisition model [16]. Segmentation-based registration techniques address intensity differences between images using anatomical structures to guide the registration. El-Rewaidy et al. utilized a non-rigid active shape model-based framework, registering T1-weighted images and computing residual motion by matching manually annotated myocardium contours. This method improves T1 map quality by addressing motion-induced errors in myocardial segments [4]. Hering et al. further advanced cardiac motion tracking by combining global semantic information from segmentation labels with local distance metrics, improving the alignment of surrounding structures [14].
@@
A key limitation of the methods above is their failure to account for the signal decay model during registration, resulting in potentially non-physical corrected T1 maps. Model-based approaches have been developed to address this, significantly enhancing T1 mapping accuracy. Xue et al. proposed a fast variational non-rigid registration framework using synthetic image estimation to correct motion in myocardial T1 mapping. Motion-free synthetic images, constructed from initial T1 estimates, guide the registration to avoid motion artifacts [31]. Van Heeswijk et al. introduced a model-based alignment of Look-Locker MRI sequences for myocardial scar quantification, using exponential curve fitting errors as the registration metric [6]. Tilborghs et al. applied a robust motion correction method for cardiac T1 and ECV mapping, integrating data-driven initialization with model-based registration within a non-rigid framework. This method iteratively estimates signal decay model parameters, generates synthetic images, and uses these for accurate registration [28]. However, their iterative nature makes these model-based approaches slower than non-iterative methods.
@@
Deep-learning approaches have also been proposed for motion correction by image registration. These networks typically learn a function that directly outputs a deformation field given an image pair. For instance, Arava et al. cascaded pairwise registration systems recursively for motion correction in T1-weighted images while minimizing the mutual information between each pair of images [1].
@@
Multi-image registration models have been employed by Hanania et al. using a mutual-information-based loss function [12]. Additionally, Gonzales et al. used a multi-scale U-Net to create distance vector fields and applied warping layers to deform feature maps from coarse to fine, which helps to align images with varying intensities by progressively refining the registration and reducing motion artifacts [8, 7]. Li et al. utilized a sparse coding technique to extract the contrast component from the images. Subsequently, they applied a self-supervised deep neural network (DNN) model for image registration [19]. Another approach by Li et al. used spatial-temporal and physical constraints to enforce low-rank and sparsity, with a deep neural network for efficient T1 mapping, denoising, and artifact reduction [18].
@@
Qin et al. used a joint learning of motion estimation and segmentation to improve registration accuracy by leveraging the unsupervised features learned during motion estimation from a large amount of unannotated data [22]. Pan et al. and Hanania et al. employed GANs to generate virtual images after style transfer at different inversion times from a single input image for image registration. The GANs reduce the sensitivity of the models to the inherent intensity differences by standardizing contrast across images, which helps to align images with varying contrasts more accurately [21, 11].
@@
Yang et al. introduced a sequential method for addressing contrast differences in images. This method initially separates intensity changes from different inversion times from the constant anatomical structure. However, it heavily relies on accurately disentangling the anatomical structure from the contrast. Additionally, the registration process only occurs between the isolated anatomical images, disregarding the signal decay model along the inversion time axis. Consequently, this approach could result in physically unrealistic deformations due to the necessity for direct integration with the signal decay model [32].
@@
These deep learning-based methods, however, do not directly consider the signal decay model, which is essential as the image contrast changes according to it. As a result, these methods may produce physically unrealistic deformations. While classical model-based approaches address these issues, their iterative nature makes them relatively slow.
@@
Previously, we introduced PCMC-T1, a method designed explicitly for free-breathing myocardial T1 mapping that incorporates physical constraints into the motion correction process [13]. This deep learning-based approach uses a supervised framework to ensure accurate T1 mapping under free-breathing conditions. However, the practical application of PCMC-T1 is limited by its reliance on large annotated datasets and extensive training, which are not readily available for every scanner and protocol.
@@
This work introduces a self-supervised approach extending our previous PCMC-T1 model. The proposed MBSS-T1 approach enables the generation of corrected T1 maps across various imaging protocols, eliminating the need for protocol-specific data collection, manual annotations, and extensive training.
@@ 3 Methods
The T1 relaxation time is described using the Inversion Recovery T1 Relaxation Equation [30]:
In this equation, represents the equilibrium magnetization, which is the magnetization of the tissue before any relaxation occurs. The parameter is the longitudinal relaxation time, indicating how quickly the magnetization returns to its equilibrium state along the longitudinal axis after inversion. The variable denotes the time at which the magnetization is measured, and is the magnetization at time , given the initial magnetization and the relaxation time .
@@
The estimation of the and parameters is typically performed by solving a least-squares optimization problem [27]. This involves minimizing the sum of the squared differences between the measured magnetization values and the values predicted by the model. The least-squares optimization problem can be formulated as follows:
@@
where represents the observed magnetization values at different time points , and is the total number of measurements.
This classical least-squares approach, as well as deep learning methods such as the one described in [9], assumes that the images are motion-free, typically achieved through prior image registration techniques. However, motion can be a significant factor in practice, especially in cardiac imaging. To account for motion, the optimization problem must be extended to include deformation fields, which can be introduced as additional variables in the least-squares problem.
The optimization problem with motion can be formulated as follows:
@@
In this extended formulation, and are the estimated relaxation and magnetization parameters, while represents the estimated deformation fields that account for motion. Specifically, denotes the -th deformation field, with . corresponds to the -th original image, and is the deformed image, reflecting the influence of motion on the observed magnetization values.
Optimizing this equation directly is challenging due to the large number of unknowns linked to the T1 relaxation model and the accompanying deformation fields. This approach can be solved iteratively using classical methods such as those described in previous studies [31, 28], but these methods tend to be slow and prone to getting stuck in local minima.
@@
To overcome these limitations, we reframe the optimization task as estimating the weights of a deep neural network () that simultaneously predicts the T1 model parameters and spatial transformations from the observed magnetization signals and corresponding time points. This is achieved using two encoder-decoder modules that operate concurrently: one for predicting the T1 parameters (parametric mapping component) and one for predicting the deformation fields (motion component). The objective function to be minimized is expressed as:
In this formulation, represents the output of the parametric mapping component, which includes the deformation fields. Specifically, corresponds to the -th deformation field. Meanwhile, denotes the output of the motion component, which includes both the T1 map () and the M0 map ().

### v3

# MBSS-T1: Model-Based Subject-Specific Self-Supervised Motion Correction for Robust Cardiac T1 Mapping

@@ Abstract
Cardiac T1 mapping is a valuable quantitative MRI technique for diagnosing diffuse myocardial diseases. Traditional methods, relying on breath-hold sequences and cardiac triggering based on an ECG signal, face challenges with patient compliance, limiting their effectiveness. Image registration can enable motion-robust cardiac T1 mapping, but inherent intensity differences between time points pose a challenge. We present MBSS-T1, a subject-specific self-supervised model for motion correction in cardiac T1 mapping. Physical constraints, implemented through a loss function comparing synthesized and motion-corrected images, enforce signal decay behavior, while anatomical constraints, applied via a Dice loss, ensure realistic deformations. The unique combination of these constraints results in motion-robust cardiac T1 mapping along the longitudinal relaxation axis. In a 5-fold experiment on a public dataset of 210 patients (STONE sequence) and an internal dataset of 19 patients (MOLLI sequence), MBSS-T1 outperformed baseline deep-learning registration methods. It achieved superior model fitting quality (: 0.975 vs. 0.941, 0.946 for STONE; 0.987 vs. 0.982, 0.965 for MOLLI free-breathing; 0.994 vs. 0.993, 0.991 for MOLLI breath-hold), anatomical alignment (Dice: 0.89 vs. 0.84, 0.88 for STONE; 0.963 vs. 0.919, 0.851 for MOLLI free-breathing; 0.954 vs. 0.924, 0.871 for MOLLI breath-hold), and visual quality (4.33 vs. 3.38, 3.66 for STONE; 4.1 vs. 3.5, 3.28 for MOLLI free-breathing; 3.79 vs. 3.15, 2.84 for MOLLI breath-hold). MBSS-T1 enables motion-robust T1 mapping for broader patient populations, overcoming challenges such as suboptimal compliance, and facilitates free-breathing cardiac T1 mapping without requiring large annotated datasets. Our code is available at https://github.com/TechnionComputationalMRILab/MBSS-T1.
keywords:
T1 Mapping, Motion Correction, Myocardial Imaging, Deep Learning, Model-based Deep Learning, Free-breathing MRI1 Introduction
@@
Quantitative cardiac T1 mapping is an advanced MRI technique designed to precisely measure intrinsic longitudinal relaxation time in myocardial tissue [32]. This method is increasingly recognized as essential for assessing diffuse myocardial diseases such as inflammation, fibrosis, hypertrophy, and infiltration [32, 35]. “Native” cardiac T1 mapping, performed without a paramagnetic contrast agent, effectively detects myocardial conditions, including edema, iron overload, infarcts, and scarring [30]. Furthermore, by allowing for the objective quantification of tissue characteristics, cardiac T1 mapping enables longitudinal tracking of changes in the myocardium. This feature is precious for clinical trials, where monitoring the efficacy of interventions over time is critical [30].
@@
Figure 1 presents the process of generating accurate T1 maps. It begins with capturing a series of images at different time points. Each image captures the myocardium at a distinct point during the longitudinal relaxation. These images are then analyzed to calculate T1 relaxation times through signal model fitting, resulting in detailed T1 maps that reflect the myocardial tissue properties. However, heart motion, respiration, and spontaneous patient movements can cause significant distortions in T1 maps, affecting their reliability and clinical usefulness and potentially leading to incorrect diagnoses [34]. Cardiac triggering based on an ECG signal is a well-established method for accounting for cardiac motion. Breath-hold sequences, such as the Modified Look-Locker Inversion Recovery (MOLLI) sequence and its variants [29], are often used to minimize respiratory motion artifacts. However, requiring patients to hold their breath presents practical challenges, as not all patients can fully adhere to these instructions. This breath-hold acquisition further limits the number of images that can be captured and makes cardiac T1 mapping unsuitable for patients who cannot tolerate breath-holding [29].
@@
Reducing motion artifacts and enabling free-breathing cardiac T1 mapping can be accomplished through image registration to align single-shot images taken at different time points. However, achieving precise registration is challenging due to inherent complexities in the image data. These complexities include contrast inversion, partial volume effects, and signal nulling for images captured near the zero crossing of the T1 relaxation curve (Figure 1).
@@
Despite advancements in classical and deep learning-based methods for motion correction and cardiac T1 mapping, significant limitations persist. Classical model-based approaches, while effective, are often slow due to their iterative nature [34, 38]. On the other hand, deep-learning-based methods rely heavily on extensively annotated datasets, which are not readily available for all scanners and protocols [10, 39]. Furthermore, these methods typically use either pairwise loss functions based on metrics such as mutual information to guide the registration process [1] or groupwise approaches such as the principal component analysis (PCA), which ignore the signal decay relaxation model [41]. This omission may result in physically unlikely deformations due to changes in image contrast. These limitations hinder the practical applicability and robustness of current methods.
@@
This work presents MBSS-T1, a subject-specific self-supervised, physically, and anatomically constrained deep learning model for simultaneous motion correction and cardiac T1 mapping from either breath-hold or free-breathing acquisitions. Our method integrates physical and anatomical insights to correct motion in cardiac T1 mapping. The main innovation of our approach lies in embedding the signal model directly into the network architecture within a subject-specific self-supervised framework. Additionally, we leverage a segmentation network pre-trained on the STONE dataset [4, 5] to generate segmentations that guide the registration process, encouraging anatomically accurate alignment of the deformed images to a reference frame. This approach enables the generation of motion-robust cardiac T1 maps across various imaging protocols, eliminating the need for protocol-specific data collection, manual annotations, and extensive training. This work extends our previous supervised PCMC-T1 model [15], which required a large annotated dataset and extensive training, thereby limiting its generalization capacity across different acquisition settings, such as sequence type and the number of frames acquired.
@@ 2 Related Work
Previous works aimed to address the challenge of motion-robust cardiac T1 mapping can be divided into classical approaches and more recent deep-learning-based approaches. Classical approaches have primarily relied on model-based techniques and iterative algorithms to mitigate the effects of intensity variations between the different frames. More recently, deep learning-based methods have emerged, leveraging large datasets and neural network architectures to enhance motion correction and cardiac T1 mapping accuracy. In the following, we review these works.
@@
Zhang et al. addressed intensity variations in their pairwise non-rigid elastic registration framework for cardiac T1 mapping by employing the normalized gradient fields difference as their similarity metric [40]. Similarly, Roujol et al. presented an adaptive registration method for varying contrast-weighted images. This approach uses a local non-rigid registration framework based on optical flow to simultaneously estimate motion fields and intensity variations, effectively managing significant contrast differences. Additionally, the method incorporates a regularization term to constrain the deformation field through automatic feature tracking [28].
@@
Guyader et al. introduced a groupwise image registration technique for quantitative MRI that utilizes a dissimilarity metric derived from an approximated form of total correlation. This method addresses the challenge of aligning multiple images simultaneously without relying on a reference image, thereby reducing registration bias and improving the consistency of quantitative measurements [12]. Tao et al. introduced a robust motion correction method for myocardial T1 and extracellular volume mapping using a PCA-based groupwise image registration technique, which registers all MOLLI frames simultaneously with a total-correlation-based metric [31]. Similarly, Huizinga et al. developed a PCA-based groupwise image registration technique for quantitative MRI (qMRI), including cardiac T1 mapping, to correct misalignment without a reference image, thus avoiding registration bias. Their method registers all images to a mean space by minimizing a PCA-derived cost function, assuming that a low-dimensional signal model can represent intensity changes in qMRI without needing details of the specific acquisition model [18].
Segmentation-based registration techniques address intensity differences between images using anatomical structures to guide the registration. El-Rewaidy et al. utilized a non-rigid active shape model-based framework, registering T1-weighted images and computing residual motion by matching manually annotated myocardium contours. This method improves T1 map quality by addressing motion-induced errors in myocardial segments [4].
@@
A key limitation of the methods above is their failure to account for the signal model during registration, resulting in potentially non-physical corrected T1 maps. Model-based approaches have been developed to address this, significantly enhancing cardiac T1 mapping accuracy. Xue et al. proposed a fast variational non-rigid registration framework using synthetic image estimation to correct motion in myocardial T1 mapping. Motion-free synthetic images, constructed from initial T1 estimates, guide the registration to avoid motion artifacts [38]. Van Heeswijk et al. introduced a model-based alignment of Look-Locker MRI sequences for myocardial scar quantification, using exponential curve fitting errors as the registration metric [8]. Tilborghs et al. applied a robust motion correction method for cardiac T1 and ECV mapping, integrating data-driven initialization with model-based registration within a non-rigid framework. This method iteratively estimates signal model parameters, generates synthetic images, and uses these for accurate registration [34]. However, their iterative nature makes these model-based approaches slower than non-iterative methods.
@@
Deep-learning approaches have also been proposed for motion correction by image registration. These networks typically learn a function that directly outputs a deformation field given an image pair. For instance, Morales et al. [24] implemented and validated a three-dimensional convolutional neural network for estimating cardiac motion from Cine MRI images. Kustner et al. [20] proposed a deep learning approach for fast and accurate non-rigid registration directly from undersampled k-space data of cardiac T1-weighted spoiled gradient echo images. Qi et al. [26] developed a supervised end-to-end deep neural network capable of simultaneously performing 3D non-rigid motion estimation and motion-compensated reconstruction from undersampled cardiac k-space data. Hering et al. [16] improved cardiac motion tracking by integrating global semantic information from segmentation labels with local distance metrics, thereby enhancing the alignment of surrounding structures. Teed et al. [33] introduced RAFT, a deep neural network employing recurrent units to estimate optical flow in videos; however, it was not specifically designed for medical imaging applications. Despite their advancements, all these methods assume relatively consistent signal appearances between the frames being registered, which limits their applicability to cardiac T1 mapping, where inherent signal variations exist across frames.
@@
Several deep-learning approaches have been proposed to address the inherent signal variations in cardiac T1 mapping. Arava et al. [1] introduced a recursive framework that cascades pairwise registration systems for motion correction in T1-weighted images, minimizing mutual information between each pair of images. Hanania et al. [14] employed multi-image registration models optimized using a mutual-information-based loss function. Gonzales et al. [10, 9] developed a multi-scale U-Net architecture to generate distance vector fields, incorporating warping layers to progressively deform feature maps from coarse to fine. This strategy enables the alignment of images with varying intensities while reducing motion artifacts through iterative refinement. Li et al. [22] applied a sparse coding technique to extract contrast components from the images, followed by a self-supervised deep neural network (DNN) for image registration. In a related work, Li et al. [21] integrated spatial-temporal and physical constraints with low-rank and sparsity regularization into a deep neural network for efficient cardiac T1 mapping, denoising, and artifact reduction.
@@
Qin et al. [27] proposed a joint learning approach for motion estimation and segmentation, leveraging unsupervised features learned from large amounts of unannotated data to improve registration accuracy. Pan et al. [25] and Hanania et al. [13] adopted generative adversarial networks (GANs) to perform style transfer, generating virtual images at different inversion times from a single input image. By standardizing contrast across images, the GAN-based methods reduce sensitivity to intensity variations and enhance the alignment of images with varying contrasts.
@@
Yang et al. [39] introduced a sequential method to address contrast differences by initially disentangling intensity changes at different inversion times from the constant anatomical structure. However, this approach critically depends on accurately separating the anatomical structure from the contrast information. Additionally, registration occurs solely between the isolated anatomical images, neglecting the signal model along the inversion time axis. As a result, this method may lead to physically unrealistic deformations due to its lack of integration with the underlying signal model.
@@
However, these approaches do not explicitly incorporate the signal model into their registration systems, despite its critical role in governing image contrast changes. Consequently, they may produce physically unrealistic deformations. In contrast, classical model-based methods account for these issues, but their iterative nature renders them comparatively slow.
@@
Previously, we introduced PCMC-T1, a method designed explicitly for free-breathing myocardial T1 mapping that incorporates physical constraints into the motion correction process [15]. This deep learning-based approach uses a supervised framework to encourage physically realistic cardiac T1 mapping under free-breathing conditions. However, the practical application of PCMC-T1 is limited by its reliance on large annotated datasets and extensive training, which are not readily available for every scanner and protocol.
@@
This work introduces MBSS-T1, a subject-specific self-supervised approach extending our previous PCMC-T1 model. The proposed MBSS-T1 approach enables the generation of motion robust T1 maps across various imaging protocols, eliminating the need for protocol-specific data collection, manual annotations, and extensive training.
@@ 3 Methods
The T1 recovery time can be modeled using distinct equations tailored to the specific sequences employed for signal measurement. The overall signal recovery for inversion recovery and Look-Locker-based measurements can generally be described as follows:
where is the magnitude image reconstructed from the multi-coil raw data obtained at time .
For the modified Look-Locker inversion recovery (MOLLI) sequence, and are fitting parameters related to the equilibrium magnetization and type of preparation, is the time after the preparation (i.e. inversion), and is the apparent which needs to be further corrected to get the actual tissue using the following equation [32]:
For the slice-interleaved T1 (STONE) sequence, the two-parameter signal recovery model (i.e. assuming a perfect efficiency of the inversion pulses) can be derived from Eq. 1 as follows: can be substituted by the equilibrium magnetization which is the magnetization of the tissue before any preparation occurs, by , and by the actual to get the proper signal recovery equation for this sequence [37]:
@@
The estimation of the signal recovery model parameters (i.e. for the MOLLI sequence and for the STONE sequence) is typically performed by solving a least-squares optimization problem [32]. This involves minimizing the sum of the squared differences between the measured magnitude images and the values predicted by the model as follows:
@@
where represents the observed magnitude images at different time points , is the total number of measurements, are the model parameters.
This classical least-squares approach, as well as deep learning methods such as the one described in [11], assumes no motion between the different images, typically achieved through prior image registration techniques. However, achieving precise registration is particularly challenging due to the inherent complexities of the image data, such as contrast inversion, partial volume effects, and signal nulling for images acquired near the zero-crossing of the T1 relaxation curve. Furthermore, sequential registration and model fitting methods may converge to local minima, leading to pairwise registrations that fail to exhibit physically accurate signal recovery. To explicitly address motion during model fitting, the optimization problem must be expanded to incorporate deformation fields, which can be treated as additional variables within the least-squares framework as follows:
@@
In this extended formulation, are the estimated signal recovery parameters (Eq. 1, or Eq. 3 according to the sequence used to acquire the data), while represents the estimated deformation fields that account for motion. Specifically, denotes the -th deformation field between the -th image and the reference image , with and corresponds to the -th original magnitude image, and is the deformed image, reflecting the influence of motion on the observed magnetization values. In case is the reference image itself, is the identity deformation field.
Direct optimization of this equation is challenging due to the high dimensionality of the problem, stemming from the numerous unknowns associated with the signal recovery model and the deformation fields. While iterative solutions using classical methods, such as those outlined in [38, 34], are feasible, they are often computationally expensive and susceptible to convergence at local minima.
@@
To overcome these limitations, we reframe the optimization task as estimating the weights of a deep neural network () that simultaneously predicts the signal recovery model parameters and spatial transformations from the observed magnitude images at the corresponding time points. This is achieved using two encoder-decoder modules that operate concurrently: one for predicting the signal recovery model parameters (parametric mapping component) and one for predicting the deformation fields (motion component). The objective function to be minimized is expressed as:
where represents the output of the parametric mapping component, are the magnitude images synthesized from the estimated parameters at the time using the signal recovery model (Eq. 1, or Eq. 3 according to the sequence used to acquire the data), and denotes the output of the motion component, which includes the predicted deformation fields between the images and the reference image .